OpenStudy (anonymous):

find an equation for the tangent line to the graph of: f(x)=1/2sec(x)-cos(x) at the point (1/3p, f(1/3p)) [p means pi]

OpenStudy (helder_edwin):

first find the derivative of the function.

OpenStudy (anonymous):

okay I got f'(x)=1/2tanxsecx+sinx what do I do now?

OpenStudy (helder_edwin):

wait. is it \[ \large f(x)=\frac{1}{2\sec x}-\cos x \] ??

OpenStudy (anonymous):

no (1/2)secx-cosx my b

OpenStudy (helder_edwin):

\[ \large f(x)=\frac{1}{2}\sec x-\cos x \] then \[ \large f'(x)=\frac{1}{2}\sec x\tan x+\sin x \] right?

OpenStudy (anonymous):

yes so where do I go from there?

OpenStudy (helder_edwin):

first compute \[ \large f(1/3\pi)= \]

OpenStudy (anonymous):

is it 1/2?

OpenStudy (helder_edwin):

\[ \large f(1/3\pi)=1/2\sec(1/3\pi)-\cos(1/3\pi)=1/2 \]

OpenStudy (helder_edwin):

then u have the point \((1/3\pi,1/2)\) the tangent line has the form \[ \large y-f(1/3\pi)=f'(1/3\pi)(x-1/3\pi) \] \[ \large y-1/2=f'(1/3\pi)(x-1/3\pi) \]

OpenStudy (anonymous):

so \[fprime(1/3\pi)=\sqrt{3}+(\sqrt{3}+4)\] ?

OpenStudy (helder_edwin):

u have to compute \[ \large f'(1/3\pi)= \]

OpenStudy (anonymous):

I did I was just wondering if I did it right

OpenStudy (helder_edwin):

i got 2.5980

OpenStudy (anonymous):

okay yeah I have to keep it in radical form but I think those answers are the same thing.

OpenStudy (helder_edwin):

\[ \large f'(1/3\pi)=1/2\sec(1/3\pi)\tan(1/3\pi)+\sin(1/3\pi) \] \[ \large =1/2\cdot 2\cdot\sqrt{3}+\frac{\sqrt{3}}{2}=3/2\sqrt{3} \]

OpenStudy (anonymous):

I don't see where you got that answer

OpenStudy (helder_edwin):

\[ \large \sin1/3\pi=\sqrt{3}/2 \] and \[ \large \cos1/3\pi=1/2 \] isn't it??

OpenStudy (anonymous):

yes but I'm not seeing how you got \[3/2\sqrt{3}\]

OpenStudy (helder_edwin):

well \[ \large =\sqrt{3}+\frac{\sqrt{3}}{2}=\sqrt{3}(1+1/2)=3/2\cdot\sqrt{3} \]

OpenStudy (anonymous):

ohhh ok got it. so the final answer should be \[2y-3\sqrt{3}x=1-\sqrt{3}\pi\]

OpenStudy (helder_edwin):

yes

OpenStudy (anonymous):

ok thanks! do you have time to help me answer one more problem?

OpenStudy (helder_edwin):

yes why not i hope it not too long

OpenStudy (anonymous):

alright thanks. here it is: a particle moves along a line so that at a time, t, where \[0\le t \le \pi,\] its position is given by \[s(t)=-4\cos t-(t^2/2)+10\]. find the velocity of the particle when its acceleration is zero.

OpenStudy (helder_edwin):

velocity if s' and acceleration is s'' right?

OpenStudy (helder_edwin):

*is not if

OpenStudy (anonymous):

yes that's right. how would I take the derivative of that equation? its kinda intimidating

OpenStudy (helder_edwin):

then compute \(s'(t)\) and \(s''(t)\): \[ \large s'(t)=4\sin t-t \] and \[ \large s''(t)=4\cos t-1 \]

OpenStudy (helder_edwin):

intimidating?? the one u did first was way harder

OpenStudy (helder_edwin):

now solve s''(x)=0

OpenStudy (anonymous):

both are intimidating. yeah I found this help site off of a website that is for k-8th grade but I'm taking ap calc. okay thanks that helped a lot. so when you plug o in for s"(x) then you get 1.3181. then you plug that t value into the velocity equation. right?

OpenStudy (helder_edwin):

yes

OpenStudy (helder_edwin):

what is K8th??

OpenStudy (anonymous):

kindergarten thru 8th grade

OpenStudy (helder_edwin):

so u r like 15??

OpenStudy (helder_edwin):

keep it up!!

OpenStudy (anonymous):

haha thanks!

OpenStudy (helder_edwin):

u r welcome glad to help